joe needs 4 pieces of lumber each 6 and 7/8 inches long . how much lumber does joe need
Answer by tutorcecilia(2152)

4 pieces x 6 and 7/8 inches = 4 X 6.875 = 27.5 inches = 27 1/2 inches
.
[Rewrite the fraction as a decimal: 7/8 = 7 divided by 8 = .875]
[Multiply 4 by 6.875]
.
or
4 (6+7/8) = 24+3.5 = 27.5 inches